OpenStudy (anonymous):

the volume of a rectangular box is limited to 1100 cu. in. if the height is 8 inches and the width is 10 inches how long can the box be ?

OpenStudy (anonymous):

\[x*8*10 = 1100\] is the equation you have set up.

OpenStudy (anonymous):

Simply solve for x.

OpenStudy (anonymous):

Are you still stuck?

OpenStudy (anonymous):

yea i really forget how to do this

OpenStudy (anonymous):

In order to solve for x, you must multiply both sides of the equation by 1/8 and 1/10 (same as dividing by 8 and dividing by 10)

OpenStudy (anonymous):

1/10 * 1/8 * x*8*10=1100*1/10 * 1/8

OpenStudy (anonymous):

The 8 and 10 cancel out on the left side of the equation, leaving you with:\[x=1100*\frac{1}{8}*\frac{1}{10}\]